# SymMaP: Improving Computational Efficiency in Linear Solvers through Symbolic Preconditioning

### v1

@@ A.2 Neural Networks for Matrix Preconditioning
Recent studies have explored the use of neural networks to improve matrix preconditioning techniques. [21, 42, 57, 67, 65, 12, 41, 36, 37, 38, 48, 28, 39, 33, 22, 43] demonstrate the effectiveness of neural networks in refining multigrid preconditioning algorithms, thus streamlining the computational process. [19] utilized Convolutional Neural Networks (CNNs) for the optimization of block Jacobi preconditioning algorithms, while [56] developed corresponding Incomplete Lower-Upper Decomposition (ILU) preconditioning algorithms leveraging machine learning insights. Although these algorithms achieved impressive results, they still face challenges such as limited interpretability and reduced computational efficiency when deployed in pure CPU environments. This paper attempts to address these issues by incorporating symbolic discovery into the framework.
@@ B.1 Overview of Matrix Preconditioning Methods
  • Jacobi Method: The Jacobi preconditioner utilizes only the diagonal elements of a matrix to precondition a linear system. By approximating the inverse of the diagonal matrix, this method is computationally simple and effective for systems with strong diagonal dominance. However, its convergence rate can be slow, and its performance diminishes for poorly conditioned or weakly diagonally dominant matrices. The Jacobi method is typically used as a baseline for comparison with more sophisticated preconditioners [53].
@@
  • Gauss-Seidel (GS) Method: The Gauss-Seidel preconditioner improves upon the Jacobi method by considering both the lower triangular and diagonal parts of the matrix in a sequential manner. Unlike the Jacobi method, which updates all variables simultaneously, the GS method updates each variable in sequence using the most recent values. This leads to faster convergence, especially for diagonally dominant matrices. However, the GS method can still struggle with poorly conditioned systems, and its forward-only approach can limit performance in some applications [53].
@@
  • Successive Over-Relaxation (SOR): The SOR method builds on the Gauss-Seidel method by introducing a relaxation factor $\omega$ to accelerate convergence. This factor allows for over-relaxation ($\omega>1$) or under-relaxation ($\omega<1$), tuning the method for faster performance on certain types of problems. SOR can significantly reduce the number of iterations needed for convergence compared to both the Jacobi and GS methods, but choosing the optimal relaxation factor is problem-dependent [68].
@@
  • Symmetric Successive Over-Relaxation (SSOR): SSOR is a symmetric version of the SOR method, where relaxation is applied in both forward and backward sweeps of the matrix. This bidirectional process improves stability and is well-suited for use with iterative solvers like the conjugate gradient method, which requires symmetric preconditioners. SSOR’s symmetry ensures that the preconditioner maintains the properties needed for efficient and stable convergence, making it a popular choice for symmetric positive-definite systems [18].
@@
  • Algebraic Multigrid (AMG): AMG is an advanced preconditioning technique designed to handle large, sparse systems of linear equations, especially those arising from the discretization of partial differential equations. Unlike traditional methods, AMG operates on multiple levels of the matrix structure, coarsening the matrix to form a hierarchy of smaller systems that are easier to solve. Solutions on the coarser grids are then interpolated back to the finer grids. This multilevel approach makes AMG highly efficient for large-scale problems, as it can dramatically reduce the number of iterations needed to achieve convergence. AMG is often used in combination with methods like SSOR or Gauss-Seidel as a smoother on each grid level, and it is particularly effective in cases where the problem exhibits a multiscale nature [51].
@@
Relationship Among Jacobi, GS, and SOR Methods: The Jacobi method is the simplest of the three, using only diagonal information. The GS method improves upon the Jacobi method by using both diagonal and lower triangular matrix elements to achieve faster convergence. SOR further refines the GS method by introducing a relaxation factor to optimize the update process. Both the GS and SOR methods can be seen as iterative improvements on the Jacobi method, with SOR offering a more flexible and potentially faster alternative by adjusting the relaxation factor. SSOR extends SOR symmetrically, making it suitable for use in more advanced iterative solvers like the conjugate gradient method [53, 18].
@@ B.2 Parameters in Matrix Preconditioning
The choice of preconditioning parameters significantly influences the effectiveness of the preconditioning process, especially in the iterative solving of linear systems [7]. Below, we discuss three specific preconditioning techniques—SOR, SSOR, and AMG—focusing particularly on how their key parameters affect the preconditioning results.
@@ B.2.1 Relaxation Factor $\omega$ in SOR and SSOR Methods
where $\bm{D}$, $\bm{L}$, and $\bm{U}$ are the diagonal, strictly lower triangular, and strictly upper triangular parts of the matrix $\bm{A}$, respectively [18].
@@
where $\bm{M}_{\text{SSOR}}$ constitutes the preconditioner, and $\bm{D}$, $\bm{L}$, $\bm{U}$, and $\omega$ are defined similarly to their roles in the SOR method. This symmetrical formulation enhances the stability and effectiveness of the preconditioning, particularly benefiting symmetric positive-definite matrices by optimizing the convergence properties of the iterative solver [18].
The choice of $\omega$ directly impacts the speed of convergence and the condition number of the matrix. Different problems and scenarios often require different choices of $\omega$, which typically need to be determined based on the specific properties of the problem and through numerical experimentation [18]. In the PETSc library, the default relaxation factor $\omega$ for both SOR and SSOR is set to 1, at which point SOR degenerates to GS preconditioning.
@@ B.2.2 Threshold Parameters $\theta_{T}$ in AMG
In the AMG method, the threshold parameter $\theta_{T}$ determines whether the non-zero elements of the matrix are "strong" enough to be considered in the construction of a coarse grid during the multigrid process. This parameter is crucial for establishing the connectivity between coarse and fine grids in the hierarchical multilevel structure [51].
@@
Different values of $\theta_{T}$ directly influence the condition number of the preconditioned matrix. Selecting the appropriate threshold parameter typically involves considering the specific structure and features of the problem, and adjustments are made through experimental fine-tuning to achieve the optimal balance [61]. In the PETSc library, the default threshold parameter $\theta_{T}$ is set to 0.
@@ 1. Darcy Flow Problem
We consider two-dimensional Darcy flows, which can be described by the following equation [35, 49, 27, 40]:
@@
In our experiment, $K(x,y)$ is generated using truncated Chebyshev polynomials. We convert the Darcy flow problem into a system of linear equations using the central difference scheme of Finite Difference Methods (FDM) [32]. The coefficients of the Chebyshev polynomials serve as input features for our symbolic discovery framework.
@@ 2. Second-order Elliptic Partial Differential Equation
We consider general two-dimensional second-order elliptic partial differential equations, which are frequently described by the following generic form [14, 6]:
@@ 3. Biharmonic Equation
We consider the biharmonic equation, a fourth-order elliptic equation, defined on a domain $\Omega\subset\mathbb{R}^{2}$. The equation is expressed as follows [9, 17, 5]:
@@
In our experiments, we construct the dataset by varying the solution domain $\Omega=[0,a]\times[0,b]$. We utilize the discontinuous Galerkin finite element method from the FEniCS library to transform this problem into a system of linear equations [5]. The parameters $a,b$ of the domain serve as input features for our symbolic discovery framework.
@@ 4. Poisson Equation
We consider a two-dimensional Poisson equation, which can be described by the following equation [65, 24, 70]:
@@
In our experiments, we address the Poisson equation within a square domain , where both the boundary conditions on all four sides and the source term $f$ on the equation’s left-hand side are generated using third-order truncated Chebyshev polynomials. The finite difference method with a central difference scheme is employed to discretize the equation into a linear system. The Chebyshev coefficients serve as parameters for our symbolic discovery framework [13].
@@ 5. Thermal Problem
We consider a two-dimensional thermal steady state equation, which can be described by the following equation [65, 55, 26]:
@@ 6. Lippmann-Schwinger Equation
We consider the 2D Lippmann-Schwinger integral equation describing quantum scattering processes [63, 2]:
@@ 7. Markov chain (Boltzmann-Distributed Brownian Motion)
We consider a Markov chain optimization problem governed by the steady-state 1D Fokker-Planck equation with a Chebyshev-parameterized potential energy field [45, 64]:
@@ E.6 Statistical Insights from the Second-Order Elliptic Equation Experiments
Dataset Matrix Size Configuration None SSOR AMG AMG+SSOR ICC [41] [41]+SSOR [41]+ICC [41]+AMG [21] [21]+SSOR [34] [41]+[34] [21]+[34] [23] [41]+[23] [21]+[23] [60] [41]+[60] [21]+[60] [24] SymMaP(SSOR) [41]+SymMaP(SSOR) [21]+SymMaP(SSOR)
@@
Dataset Matrix Size Configuration None SOR AMG AMG+SOR ILU [41] [41]+SOR [41]+ILU [41]+AMG [21] [21]+SOR [34] [41]+[34] [21]+[34] [23] [41]+[23] [21]+[23] [60] [41]+[60] [21]+[60] [24] SymMaP(SOR) [41]+SymMaP(SOR) [21]+SymMaP(SOR)
@@ E.7 Comparative Experiments with Learning-Based Approaches
    Generality: SymMaP is highly adaptable to diverse preconditioners and linear solvers. In contrast, existing learning-based methods either: (a) combine with specific preconditioners (e.g., [41], [21]), or (b) target narrow use cases (e.g., [60], [23] for ICC-preconditioned CG; [34] exclusively for CG).
@@ E.7.1 Notes on These Methods
    Methods Enhancing Linear Solvers: These approaches add optimization modules or modify the iteration process (e.g., [41], [21]). These methods target different components than SymMaP and could potentially be combined with our approach for further improvements.
@@
    Methods Predicting Preconditioning Matrices: These approaches use neural networks to predict preconditioning matrices (e.g., [60], [34], [24]).
@@
  • Combinations: Sequential application of methods (e.g., ‘AMG+?‘ or ‘[41]+?‘).
  • ×: Method incompatible (e.g., [34], [23], [24], [60] only work for symmetric matrices/CG) or excessively slow (e.g., [24] >100s for Elliptic PDE).
@@
  • [41]: Accelerates Krylov solvers via recycling (distinct from preconditioning); combinable with other methods.
  • [21]: An AMG variant; allows nested preconditioning but conflicts with learning-based methods ([34], [23], [60]).
  • [24]: A fixed-point iteration (non-Krylov); computationally prohibitive for large sparse systems.
@@ E.7.2 Results
As shown in Tables 15 and 15, across all experiments, the combination of "[41]+SymMaP" consistently achieved the shortest computation times. When evaluating standalone algorithms (without combinations), SymMaP alone demonstrated the fastest performance in all test cases. The performance advantage of SymMaP was particularly pronounced in CPU-only environments. SymMaP generalizes to any parameterized preconditioner, unlike other methods (e.g., [23] for ICC, [21] for AMG). These results conclusively demonstrate the superior performance characteristics of the SymMaP approach.

### v2

@@ A.2 Neural Networks for Matrix Preconditioning
Recent studies have explored the use of neural networks to improve matrix preconditioning techniques. [greenfeld2019learning, luz2020learning, taghibakhshi2021optimization, Yang2025gnnspai, wang2024accelerating, dong2024accelerating, luo2024neural, liu2023promotinggeneralizationexactsolvers, liu2025apollomilp, liu2024milpstudio, 2025rome, kuang2024rethinking, liu2024deep, li2024machine, optitree, lv2025exploiting] demonstrate the effectiveness of neural networks in refining multigrid preconditioning algorithms, thus streamlining the computational process. [gotz2018machine] utilized Convolutional Neural Networks (CNNs) for the optimization of block Jacobi preconditioning algorithms, while [stanaityte2020ilu] developed corresponding Incomplete Lower-Upper Decomposition (ILU) preconditioning algorithms leveraging machine learning insights. Although these algorithms achieved impressive results, they still face challenges such as limited interpretability and reduced computational efficiency when deployed in pure CPU environments. This paper attempts to address these issues by incorporating symbolic discovery into the framework.
@@ B.1 Overview of Matrix Preconditioning Methods
  • Jacobi Method: The Jacobi preconditioner utilizes only the diagonal elements of a matrix to precondition a linear system. By approximating the inverse of the diagonal matrix, this method is computationally simple and effective for systems with strong diagonal dominance. However, its convergence rate can be slow, and its performance diminishes for poorly conditioned or weakly diagonally dominant matrices. The Jacobi method is typically used as a baseline for comparison with more sophisticated preconditioners [saad2003iterative].
@@
  • Gauss-Seidel (GS) Method: The Gauss-Seidel preconditioner improves upon the Jacobi method by considering both the lower triangular and diagonal parts of the matrix in a sequential manner. Unlike the Jacobi method, which updates all variables simultaneously, the GS method updates each variable in sequence using the most recent values. This leads to faster convergence, especially for diagonally dominant matrices. However, the GS method can still struggle with poorly conditioned systems, and its forward-only approach can limit performance in some applications [saad2003iterative].
@@
  • Successive Over-Relaxation (SOR): The SOR method builds on the Gauss-Seidel method by introducing a relaxation factor $\omega$ to accelerate convergence. This factor allows for over-relaxation ($\omega>1$) or under-relaxation ($\omega<1$), tuning the method for faster performance on certain types of problems. SOR can significantly reduce the number of iterations needed for convergence compared to both the Jacobi and GS methods, but choosing the optimal relaxation factor is problem-dependent [young1954iterative].
@@
  • Symmetric Successive Over-Relaxation (SSOR): SSOR is a symmetric version of the SOR method, where relaxation is applied in both forward and backward sweeps of the matrix. This bidirectional process improves stability and is well-suited for use with iterative solvers like the conjugate gradient method, which requires symmetric preconditioners. SSOR’s symmetry ensures that the preconditioner maintains the properties needed for efficient and stable convergence, making it a popular choice for symmetric positive-definite systems [golub2013matrix].
@@
  • Algebraic Multigrid (AMG): AMG is an advanced preconditioning technique designed to handle large, sparse systems of linear equations, especially those arising from the discretization of partial differential equations. Unlike traditional methods, AMG operates on multiple levels of the matrix structure, coarsening the matrix to form a hierarchy of smaller systems that are easier to solve. Solutions on the coarser grids are then interpolated back to the finer grids. This multilevel approach makes AMG highly efficient for large-scale problems, as it can dramatically reduce the number of iterations needed to achieve convergence. AMG is often used in combination with methods like SSOR or Gauss-Seidel as a smoother on each grid level, and it is particularly effective in cases where the problem exhibits a multiscale nature [ruge1987algebraic].
@@
Relationship Among Jacobi, GS, and SOR Methods: The Jacobi method is the simplest of the three, using only diagonal information. The GS method improves upon the Jacobi method by using both diagonal and lower triangular matrix elements to achieve faster convergence. SOR further refines the GS method by introducing a relaxation factor to optimize the update process. Both the GS and SOR methods can be seen as iterative improvements on the Jacobi method, with SOR offering a more flexible and potentially faster alternative by adjusting the relaxation factor. SSOR extends SOR symmetrically, making it suitable for use in more advanced iterative solvers like the conjugate gradient method [saad2003iterative, golub2013matrix].
@@ B.2 Parameters in Matrix Preconditioning
The choice of preconditioning parameters significantly influences the effectiveness of the preconditioning process, especially in the iterative solving of linear systems [chen2005matrix]. Below, we discuss three specific preconditioning techniques—SOR, SSOR, and AMG—focusing particularly on how their key parameters affect the preconditioning results.
@@ B.2.1 Relaxation Factor $\omega$ in SOR and SSOR Methods
where $\bm{D}$, $\bm{L}$, and $\bm{U}$ are the diagonal, strictly lower triangular, and strictly upper triangular parts of the matrix $\bm{A}$, respectively [golub2013matrix].
@@
where $\bm{M}_{\text{SSOR}}$ constitutes the preconditioner, and $\bm{D}$, $\bm{L}$, $\bm{U}$, and $\omega$ are defined similarly to their roles in the SOR method. This symmetrical formulation enhances the stability and effectiveness of the preconditioning, particularly benefiting symmetric positive-definite matrices by optimizing the convergence properties of the iterative solver [golub2013matrix].
The choice of $\omega$ directly impacts the speed of convergence and the condition number of the matrix. Different problems and scenarios often require different choices of $\omega$, which typically need to be determined based on the specific properties of the problem and through numerical experimentation [golub2013matrix]. In the PETSc library, the default relaxation factor $\omega$ for both SOR and SSOR is set to 1, at which point SOR degenerates to GS preconditioning.
@@ B.2.2 Threshold Parameters $\theta_{T}$ in AMG
In the AMG method, the threshold parameter $\theta_{T}$ determines whether the non-zero elements of the matrix are "strong" enough to be considered in the construction of a coarse grid during the multigrid process. This parameter is crucial for establishing the connectivity between coarse and fine grids in the hierarchical multilevel structure [ruge1987algebraic].
@@
Different values of $\theta_{T}$ directly influence the condition number of the preconditioned matrix. Selecting the appropriate threshold parameter typically involves considering the specific structure and features of the problem, and adjustments are made through experimental fine-tuning to achieve the optimal balance [trottenberg2000multigrid]. In the PETSc library, the default threshold parameter $\theta_{T}$ is set to 0.
@@ 1. Darcy Flow Problem
We consider two-dimensional Darcy flows, which can be described by the following equation [li2020fourier, rahman2022u, kovachki2021neural, lu2022comprehensive]:
@@
In our experiment, $K(x,y)$ is generated using truncated Chebyshev polynomials. We convert the Darcy flow problem into a system of linear equations using the central difference scheme of Finite Difference Methods (FDM) [leveque2007finite]. The coefficients of the Chebyshev polynomials serve as input features for our symbolic discovery framework.
@@ 2. Second-order Elliptic Partial Differential Equation
We consider general two-dimensional second-order elliptic partial differential equations, which are frequently described by the following generic form [evans2022partial, bers1964partial]:
@@ 3. Biharmonic Equation
We consider the biharmonic equation, a fourth-order elliptic equation, defined on a domain $\Omega\subset\mathbb{R}^{2}$. The equation is expressed as follows [ciarlet1974mixed, glowinski1979numerical, barrata2023dolfinx]:
@@
In our experiments, we construct the dataset by varying the solution domain $\Omega=[0,a]\times[0,b]$. We utilize the discontinuous Galerkin finite element method from the FEniCS library to transform this problem into a system of linear equations [barrata2023dolfinx]. The parameters $a,b$ of the domain serve as input features for our symbolic discovery framework.
@@ 4. Poisson Equation
We consider a two-dimensional Poisson equation, which can be described by the following equation [wang2024accelerating, hsieh2019learning, zhang2022hybrid]:
@@
In our experiments, we address the Poisson equation within a square domain , where both the boundary conditions on all four sides and the source term $f$ on the equation’s left-hand side are generated using third-order truncated Chebyshev polynomials. The finite difference method with a central difference scheme is employed to discretize the equation into a linear system. The Chebyshev coefficients serve as parameters for our symbolic discovery framework [driscoll2014chebfun].
@@ 5. Thermal Problem
We consider a two-dimensional thermal steady state equation, which can be described by the following equation [wang2024accelerating, sharma2018weakly, koric2023data]:
@@ 6. Lippmann-Schwinger Equation
We consider the 2D Lippmann-Schwinger integral equation describing quantum scattering processes [vainikko2000fast, ambikasaran2016fast]:
@@ 7. Markov chain (Boltzmann-Distributed Brownian Motion)
We consider a Markov chain optimization problem governed by the steady-state 1D Fokker-Planck equation with a Chebyshev-parameterized potential energy field [pawula1967approximation, vesely1994computational]:
@@ E.6 Statistical Insights from the Second-Order Elliptic Equation Experiments
Dataset Matrix Size Configuration None SSOR AMG AMG+SSOR ICC [luo2024neural] [luo2024neural]+SSOR [luo2024neural]+ICC [luo2024neural]+AMG [greenfeld2019learning] [greenfeld2019learning]+SSOR [li2023learning] [luo2024neural]+[li2023learning] [greenfeld2019learning]+[li2023learning] [hausner2024neural] [luo2024neural]+[hausner2024neural] [greenfeld2019learning]+[hausner2024neural] [trifonov2024learning] [luo2024neural]+[trifonov2024learning] [greenfeld2019learning]+[trifonov2024learning] [hsieh2019learning] SymMaP(SSOR) [luo2024neural]+SymMaP(SSOR) [greenfeld2019learning]+SymMaP(SSOR)
@@
Dataset Matrix Size Configuration None SOR AMG AMG+SOR ILU [luo2024neural] [luo2024neural]+SOR [luo2024neural]+ILU [luo2024neural]+AMG [greenfeld2019learning] [greenfeld2019learning]+SOR [li2023learning] [luo2024neural]+[li2023learning] [greenfeld2019learning]+[li2023learning] [hausner2024neural] [luo2024neural]+[hausner2024neural] [greenfeld2019learning]+[hausner2024neural] [trifonov2024learning] [luo2024neural]+[trifonov2024learning] [greenfeld2019learning]+[trifonov2024learning] [hsieh2019learning] SymMaP(SOR) [luo2024neural]+SymMaP(SOR) [greenfeld2019learning]+SymMaP(SOR)
@@ E.7 Comparative Experiments with Learning-Based Approaches
    Generality: SymMaP is highly adaptable to diverse preconditioners and linear solvers. In contrast, existing learning-based methods either: (a) combine with specific preconditioners (e.g., [luo2024neural], [greenfeld2019learning]), or (b) target narrow use cases (e.g., [trifonov2024learning], [hausner2024neural] for ICC-preconditioned CG; [li2023learning] exclusively for CG).
@@ E.7.1 Notes on These Methods
    Methods Enhancing Linear Solvers: These approaches add optimization modules or modify the iteration process (e.g., [luo2024neural], [greenfeld2019learning]). These methods target different components than SymMaP and could potentially be combined with our approach for further improvements.
@@
    Methods Predicting Preconditioning Matrices: These approaches use neural networks to predict preconditioning matrices (e.g., [trifonov2024learning], [li2023learning], [hsieh2019learning]).
@@
  • Combinations: Sequential application of methods (e.g., ‘AMG+?‘ or ‘[luo2024neural]+?‘).
  • ×: Method incompatible (e.g., [li2023learning], [hausner2024neural], [hsieh2019learning], [trifonov2024learning] only work for symmetric matrices/CG) or excessively slow (e.g., [hsieh2019learning] >100s for Elliptic PDE).
@@
  • [luo2024neural]: Accelerates Krylov solvers via recycling (distinct from preconditioning); combinable with other methods.
  • [greenfeld2019learning]: An AMG variant; allows nested preconditioning but conflicts with learning-based methods ([li2023learning], [hausner2024neural], [trifonov2024learning]).
  • [hsieh2019learning]: A fixed-point iteration (non-Krylov); computationally prohibitive for large sparse systems.
@@ E.7.2 Results
As shown in Tables 15 and 15, across all experiments, the combination of "[luo2024neural]+SymMaP" consistently achieved the shortest computation times. When evaluating standalone algorithms (without combinations), SymMaP alone demonstrated the fastest performance in all test cases. The performance advantage of SymMaP was particularly pronounced in CPU-only environments. SymMaP generalizes to any parameterized preconditioner, unlike other methods (e.g., [hausner2024neural] for ICC, [greenfeld2019learning] for AMG). These results conclusively demonstrate the superior performance characteristics of the SymMaP approach.